Part G For the 3.00 Ω resistor, calculate the current through the resistor with S open. ur answer with the appropriate units. LA 1- Value Units Submit My Answers Give UpPlease
Answer #1

for switch open, net resistance, Ropen = R2 + R3 = 6 + 3 = 9 ohm

for switch close, net resistance, Rclose = (R1R2/(R1 + R2)) + R3 = (24/10) + 3 = 5.4 ohm

total current for switch open = E/Ropen = 41/9 A

total current for switch close = E/Rclose = 41/5.4 A

A) Vab = IopenR2 = (41/9) x 6 = 82/3 = 27.333 V

B) Vab = E - IcloseR3 = 41 - (41/5.4) x 3 = 41(4/9) = 164/9 = 18.2222 V

C) current through 4 ohm resistor = 0

D) current through 4 ohm resistor = 6Iclose/(6+4) = 0.6Iclose = 0.6 x 41/5.4 = 41/9 = 4.5555 A

E) current through 6 ohm resistor = Iopen = 41/9 = 4.5555 A

F) current through 6 ohm resistor = 4Iclose/(6+4) = 0.4Iclose = 0.4 x 41/5.4 = 82/27 = 3.037 A

G) current through 3 ohm resistor = Iopen = 41/9 = 4.5555 A

H) current through 3 ohm resistor = Iclose = 41/5.4 = 7.5926 A

I) Current through R1 increase

Current through R2 decrease

Current through R3 increase